OpenStudy (anonymous):

Identify the x and y intercepts for the equation 3x + 4y = 12.

OpenStudy (saifoo.khan):

when x intercpts, y = 0 when y intercpts, x = 0

OpenStudy (anonymous):

huh?

OpenStudy (saifoo.khan):

when x intercpts, y = 0 3x + 4y = 12. 3x + 4(0) = 12 3x = 12 x = 12/3 x = 4 when y intercpts, x = 0 3x + 4y = 12. 3(0) + 4y = 12 4y = 12 y = 12/4 y = 3

OpenStudy (saifoo.khan):

So, (4,0) & (0,3)

OpenStudy (anonymous):

o.O oh

OpenStudy (saifoo.khan):

do u get it? lol.

OpenStudy (anonymous):

kinda..like substitution right?

OpenStudy (saifoo.khan):

Kind of. but not really.